OpenStudy (anonymous):

How to solve V(h)=h(h-10)(h-8) find maximum for the domain 0

OpenStudy (anonymous):

Take the derivative with respect to h, first, realize that for the product rule of 3 functions we have: \[(fgh)'=f'gh+fg'h+fgh'\] An extension of the 2 factor one. So we have: \[\frac{d V}{dh}=(1)(h-10)(h-8)+h(1)(h-8)+h(h-10)(1)=0\] \[\implies (h-10)((h-8)+h)+(h-8)h=0 \implies h^2+80-18h+h^2-10h+h^2-8h=0\] \[\implies h^2+80-18h+h^2-10h+h^2-8h=0 \] \[\implies 3h^2-36h+80=0 \implies h=\frac{36 \pm \sqrt{36^2-4*3*80}}{6} \implies h=\frac{2}{3}(9 \pm \sqrt{21})\] Both are within 0<h<10

OpenStudy (anonymous):

So then take the second derivative and test the concavity at these points to see if its a max or min